OpenStudy (anonymous):

If p varies directly with q, and p = 10 when q = 2, what is the value of p when q = 20? A. p = 40 B. p = 1 C. p = 100 D. p = 28

OpenStudy (texaschic101):

10/2 = p/20 2p = 200 p = 100

OpenStudy (anonymous):

thx u

OpenStudy (texaschic101):

yw